Thursday, 9 am MT, 11 am ET, 14 July 2016


CV-331/AOC-317/GB-bsm, 331 Hub Audio 434-817-6286


  • 17A
    • PST Issues (Dana, Rick)
    • PST Review Items (Dana, Rick)
    • PHT Requirements (Gustaaf, Barry)
    • SRP Recruitment (Mark C., Dave)

  • AOB


Attending: BarryClark, RickLively, HeidiMedlin, MarkClaussen, ClaireChandler, BryanButler, ToneyMinter, DavidFrayer, GarethHunt, DanaBalser

  • 17A
    • PST Issues. Thus far we have had two 17A proposals submitted (both VLA), and a couple of helpdesk tickets. RickLively fixed a bug in the GBT VEGAS resource where the integration time was not updated for Spectrometers 7 and 8. He also fixed a few cosmetics issues. There were about 20 udb-admin tickets; mostly institutions and user requests.
    • PST Review Items. RickLively has almost completed 2 of 6 items. He should be able to get to most of them.
    • PHT Requirements. GustaafVanMoorsel and BarryClark still need to get together and review the current list to prioritize existing tickets. We should probably fix the CSV file problems. BarryClark has started work on 17A updates to the VLA Prioritizer: adding a properties file and reducing the amount of high frequency filler. ClaireChandler suggested that we should also reduce the amount of high frequency A and B time, or set this more accurately based on our expectations. This will be more difficult to implement, but BarryClark will look into it. The GB PHT is undergoing some changes. PaulMarganian and RickLively are in contact and will make sure there are no communication issues between the two tools.
    • SRP Recruitment. DanaBalser reported that JayLockman is working on the 1 EGS vacancy. MarkClaussen reported he is working on the 1 ETP vacancy.

-- DanaBalser - 2016-07-12
Topic revision: r3 - 2016-07-14, DanaBalser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback